OpenStudy (itsmichelle29):

I give medals to helpers thanks Regan is trying to find the equation of a quadratic that has a focus of (−2, 5) and a directrix of y = 13. Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand.

OpenStudy (itsmichelle29):

can you guys help me plz

OpenStudy (anonymous):

y = -1/16(x - (-2))^2 +9.

OpenStudy (itsmichelle29):

I dont understand

OpenStudy (itsmichelle29):

Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand. Can you give me this answer

OpenStudy (misty1212):

HI!!

OpenStudy (itsmichelle29):

hi omg can you please help me

OpenStudy (misty1212):

no one likes these conic section questions, but they are not that hard really

OpenStudy (misty1212):

first off, do you know what the parabola would look like (open right, left, up or down?)

OpenStudy (itsmichelle29):

Nope

OpenStudy (misty1212):

hmm that is a problem lets plot the focus and graph the directrix

OpenStudy (itsmichelle29):

okay

OpenStudy (misty1212):



OpenStudy (misty1212):

you can see that the focus \((-2,5)\) is BELOW the directrix, which is a horizontal line that means the parabola will open DOWN

OpenStudy (itsmichelle29):

Ohhh okay

OpenStudy (misty1212):

the next thing we need to do is find the vertex do you know how to find it ?

OpenStudy (itsmichelle29):

No ='(

OpenStudy (misty1212):

it is the point half way between \((-2,5)\) and \(y=13\) what is half way between \(5\) and \(13\)?

OpenStudy (itsmichelle29):

9

OpenStudy (misty1212):

right (got logged out somehow)

OpenStudy (misty1212):

that means the vertex is \((-2,9)\)

OpenStudy (misty1212):

now we are almost done

OpenStudy (itsmichelle29):

OKAY

OpenStudy (itsmichelle29):

so now what ?

OpenStudy (misty1212):

since the vertex is \((-2,5)\) and it opens down, it will look like \[4p(y-9)=(x+2)^2\] all we need is \(p\) and to note that \(4p\) will be negative since it opens down

OpenStudy (misty1212):

oops i meant "since the vertex is \((-2,9)\) my mistake

OpenStudy (itsmichelle29):

no its -2 5

OpenStudy (misty1212):

to find \(p\) is also easy it is the distance between the focus \((-2,5)\) and the vertex \((-2,9)\)

OpenStudy (itsmichelle29):

ohhh n.m.

OpenStudy (misty1212):

how far from \(5\) to \(9\)?

OpenStudy (misty1212):

did i lose you?